OpenStudy (dakid88):

The area of the scale model of a garden is 15 square feet. The scale model is enlarged by a scale factor of 3 to create the actual garden. Which expression finds the area of the actual garden?

OpenStudy (anonymous):

i don't know what u would do

OpenStudy (dakid88):

the answer choices are \[\frac{ 15 }{ 3^{2} }\] \[\frac{ 15 }{ 3 }\] \[15\times3\] \[15\times3^{2}\]

OpenStudy (tkhunny):

Linear Scale Factor 3 Area Scale Factor 9 = 3^2 Volume Scale Factor 27 = 3^3

OpenStudy (dakid88):

those ar ethe answer choices

OpenStudy (tkhunny):

Okay, feel free to find the right one. It's on there.

OpenStudy (dakid88):

k anymore thought or help

OpenStudy (tkhunny):

I can't just give you the answer. That would be contrary to the OpenStudy Code of Conduct. You have all the information you need. Is the model larger or smaller than the actual?

OpenStudy (dakid88):

i know tht im talking bout help plz @tkhunny

OpenStudy (tkhunny):

Answer the question. Is the model larger or smaller than the actual?

OpenStudy (dakid88):

larger i think

OpenStudy (tkhunny):

Do better than that. "The scale model is enlarged by a scale factor of 3" The model is larger or smaller than the actual?

OpenStudy (dakid88):

smaller

OpenStudy (tkhunny):

Try again. "The scale model is enlarged"

OpenStudy (dakid88):

so it is larger

OpenStudy (tkhunny):

That's the spirit. So the actual garden is rather small, it would seem - although this is odd. Which of the answers produces a result that is less than 15? A.15/3^2 B.15/3 C.15×3 D.15×3^2

OpenStudy (dakid88):

i will say B.

OpenStudy (tkhunny):

A and B produce results less than 15. Which is correct?

OpenStudy (dakid88):

A.

OpenStudy (tkhunny):

Why?

OpenStudy (tkhunny):

Note: I'm worried about this problem. It is VERY unusual for a scale model of something like a garden to be larger than the garden. We MAY be going the wrong direction. If this one comes back marked incorrect, then it is a very poorly worded question. Definitely 15/9 - AS STATED..

OpenStudy (tkhunny):

My point is this. If it were a scale model of an atom, I would expect a model to be larger. A garden? Very odd. 15 sq ft isn't much. 1/9 of that? It's a bonsai garden, at best.